OpenStudy (anonymous):

how do i simplify (2x)^-3

OpenStudy (mertsj):

Use this property: \[a ^{-n}=\frac{1}{a^n}\]

OpenStudy (anonymous):

I am confused can you explain

OpenStudy (mertsj):

I will show you an example: \[(3x)^{-4}=\frac{1}{(3x)^4}=\frac{1}{81x^4}\]

OpenStudy (anonymous):

Got it! (2x)^-3 =1/(2x)^3 =1/8x^3

OpenStudy (mertsj):

Perfect!! Good job!!

OpenStudy (anonymous):

Thanks

OpenStudy (mertsj):

yw